OpenStudy (anonymous):

show all step,,,,,,One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square

OpenStudy (anonymous):

5n = n + 16 4n = 16 n = 4 side of square is 5*4 = 20 inches